# Full quantum tomography of top quark decays

### v1

@@ 5 Experimental sensitivity
Figure 1: Distribution of $LW$ and $Lb$ entanglement measures $N$ obtained from pseudo-experiments. The curve shows a Gaussian fit to the $N(\rho_{LW})$ distribution in the interval $[-3\sigma,3\sigma]$.
The result of 20000 pseudo-experiments is presented in Fig. 1. For $LW$ entanglement the distribution is Gaussian to a good approximation, and the expected statistical precision is $N(\rho_{LW})=0.512\pm 0.015$. For $Lb$ entanglement the distribution presents a small shoulder caused by the boundary condition $F_{+}\geq 0$. In this case, the mean is $\langle N(\rho_{Lb})\rangle=0.356$ and the standard deviation is $0.014$. In both cases, the measurements have a statistical precision that would allow to establish the entanglement ($N>0$) beyond the $5\sigma$ criterion, nominally 35 and 25 standard deviations.
@@ 6 Discussion
The reader may already have noticed two conspicuous features that arise in our theoretical framework. The $LW$ density operator, obtained after trace on the $b$ spin degrees of freedom, is still sensitive to the phase $e^{i\alpha}$ between amplitudes with $b$ quark helicities ±$\pm 1/2$. This is a consequence of the ambiguity in the direct determination of the density operators involving $L$, which we solve by expressing them in terms of $a_{\lambda_{1}\lambda_{2}}$. Even more striking is the fact that the $Lb$ entanglement can be established when the $b$ quark spin cannot be directly measured. This is possible due the change of basis used to write the $LWb$ density operator in terms of a minimal set of parameters $a_{\lambda_{1}\lambda_{2}}$, and the fact that $\lambda=1/2$ are suppressed.
The excellent statistical sensitivity expected for these entanglement measurements suggests that they will be limited by systematic uncertainties. For the single top polarisation, the Run 2 ATLAS measurement [18], which used a multi-dimensional template fit, has uncertainties of 17%, 2.5% and 10%, for $P_{1}$, $P_{2}$ and $P_{3}$, respectively. The measurements of the $W$ helicity fractions by the CMS Collaboration in single top production with Run 1 data [19] have systematic uncertainties below 4%. Asymmetries equivalent to $A_{\text{FB}}^{x^{\prime}}$, $A_{\text{FB}}^{y^{\prime}}$, $A_{\text{FB}}^{x^{\prime}z^{\prime}}$, $A_{\text{FB}}^{y^{\prime}z^{\prime}}$ have been measured by the ATLAS Collaboration in single top production with Run 1 data [20] with systematics of 8%, 3%, 11% and 5%, respectively. A measurement of $A_{\text{FB,EC}}^{z,z^{\prime}}$ has not been performed yet. It requires reconstruction of the $\hat{z}$ axis (spectator jet) and $\hat{z}^{\prime}$ axis ($W$ momentum). In both cases the directions are well determined, so we expect a systematic uncertainty at the 10% level, as obtained for $P_{3}$. We also remark that with high statistics, as those available with Run 2 data, the full distributions can be used for measurements, not only asymmetries. With that purpose, a deconvolution of the detector effects for measurements of the four-dimensional distribution has been proposed [21, 22], which also takes into account the correlations among measurements.
In summary, the determination of the $LW$ and $Lb$ entanglement is based on observables which, except one, have already been measured by the ATLAS or CMS Collaborations, in separate analyses. The current experimental uncertainties suggest that a dedicated analysis of Run 2 data could establish the entanglement between $L$ and the $W$ boson spin, or $L$ and the $b$ quark spin, with more than 5 standard deviations.

### v2

@@ 5 Experimental sensitivity
We estimate the statistical sensitivity for the determination of the entanglement measures for bipartite and tripartite entanglement with the data collected at the LHC Run 2, working at the partonic level. An event sample is generated with MadGraph [35], in the process $pp\to t\bar{b}j$, using NNPDF 3.1 [36] parton density functions and setting as factorisation and renormalisation scale the average transverse mass. The resulting top quark polarisation is $P_{3}=0.845$ in the spectator jet axis, with $P_{1}=0.035$, $P_{2}=0$ in orthogonal directions [37, 38]. We perform pseudo-experiments, in which (i) a subset of events is randomly chosen; (ii) the values of $P_{i}$ and $a_{\lambda_{1}\lambda_{2}}$ are determined from the set; (iii) the density operators are reconstructed using Eqs. (3) and (9); (iv) the entanglement measures are computed. For each pseudo-experiment we take a sample of 50000 events, which is the expected number of the events after selection cuts in the single top polarisation measurement performed by the ATLAS Collaboration [38] with the full Run 2 luminosity (139 fb-1).
There are two points that are worth being noted, concerning our procedure. First, the background from other processes is about the same size as the signal [38], and its statistical fluctuations will increase the signal uncertainty. However, since the stastistical uncertainties obtained with pseudo-experiments are well beyond $5\sigma$ sensitivity, this effect turns out to be irrelevant in order to establish the observation of entanglement. The second point concerns the relative phase that cannot be measured, which is set to its SM value, $\alpha=0$. Using a non-SM value increases $N(\rho_{LW})$ up to $0.013$ and $N(\rho_{Lb})$ up to $0.017$. The impact of this phase on the entanglement between the different bipartitions of $\mathcal{H}_{L}\otimes\mathcal{H}_{W}\otimes\mathcal{H}_{b}$ is also negligible: $N(\rho_{L[Wb]})$ is increased by up to $0.030$ and $N(\rho_{b[LW]})$ decreases up to $-1.4\times 10^{-3}$. Again, the effect is irrelevant in order to establish the observation of entanglement.
Refer to caption Refer to caption
Figure 1: Distribution of entanglement measures $N$ obtained from pseudo-experiments, for bipartite entaglement (left) and tripartite entanglement (right). The curves show, in the interval $[-3\sigma,3\sigma]$, a Gaussian fit to the corresponding distribution.
For bipartite entanglement, the result of 20000 pseudo-experiments is presented in the left panel of Fig. 1. For $L$-$W$ entanglement the distribution is Gaussian to a good approximation, and the expected statistical precision is $N(\rho_{LW})=0.512\pm 0.015$. For $L$-$b$ entanglement the distribution presents a small shoulder caused by the boundary condition $F_{+}\geq 0$. In this case, the mean is $\langle N(\rho_{Lb})\rangle=0.356$ and the standard deviation is $0.014$. In both cases, the measurements have a statistical precision that would allow to establish the entanglement ($N>0$) beyond the $5\sigma$ criterion, nominally 35 and 25 standard deviations.
For tripartite entanglement, the result of 40000 pseudo-experiments is presented in the right panel of Fig. 1. For $W$-$(Lb)$ and $b$-$(LW)$ the distributions are Gaussian, with $N(\rho_{W[Lb]})=0.701\pm 0.006$, $N(\rho_{b[LW]})=0.496\pm 0.004$. For $L$-$(Wb)$ the distribution has a small shoulder on the right. The mean and standard deviation are in this case $\langle N(\rho_{L[Wb]})\rangle=1.42$, $\sigma=0.03$. The statistical difference from the separability hypothesis $N=0$ is, in all cases, much larger than 5 standard deviations, and would allow to establish genuine tripartite entanglement.
The excellent statistical sensitivity expected for these entanglement measurements suggests that they will be limited by systematic uncertainties. For the single top polarisation, the Run 2 ATLAS measurement [38], which used a multi-dimensional template fit, has uncertainties of 17%, 2.5% and 10%, for $P_{1}$, $P_{2}$ and $P_{3}$, respectively. The measurements of the $W$ helicity fractions by the CMS Collaboration in single top production with 20 fb-1 of Run 1 data [39] have systematic uncertainties below 4%. Asymmetries equivalent to $A_{\text{FB}}^{x^{\prime}}$, $A_{\text{FB}}^{y^{\prime}}$, $A_{\text{FB}}^{x^{\prime}z^{\prime}}$, $A_{\text{FB}}^{y^{\prime}z^{\prime}}$ have been measured by the ATLAS Collaboration in single top production with 20 fb-1 of Run 1 data [40], finding systematics of 8%, 3%, 11% and 5%, respectively. A measurement of $A_{\text{FB,EC}}^{z,z^{\prime}}$ has not been performed yet. It requires reconstruction of the $\hat{z}$ axis (spectator jet) and $\hat{z}^{\prime}$ axis ($W$ momentum). In both cases the directions are well determined, so we expect a systematic uncertainty at the 10% level, as obtained for $P_{3}$. Taking as central values the SM predictions and using the systematic uncertainties in Refs. [38, 39, 40] as benchmark, one can calculate the variation in entanglement observables due to (uncorrelated) variations in the input observables. The results are presented in Table 1. Summing in quadrature all the different contributions, one can thus estimate the total systematic uncertainty on each entanglement measure, summarised in the last row of Table 1. (This procedure does not take into account correlations between observables, which must be calculated once a simultaneous measurement of them all is performed.)
@@
[…]
For all entanglement measures except $N(\rho_{Lb})$, the sensitivity to establish entanglement ($N>0$) remains above the $5\sigma$ level after including the estimation of systematic uncertainties.
$N(\rho_{LW})$ $N(\rho_{Lb})$ $N(\rho_{L[Wb]})$ $N(\rho_{W[Lb]})$ $N(\rho_{b[LW]})$
$P_{1}$ ${}^{+0.01}_{-0}$ ${}^{+0.003}_{-0}$ ${}^{+0.03}_{-0}$ ${}^{+0.044}_{-0.003}$ ${}^{+0.021}_{-0.003}$
$P_{2}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0.001}_{-0}$ ${}^{+0}_{-0}$
$P_{3}$ ${}^{+0.015}_{-0.024}$ ${}^{+0}_{-0.013}$ ${}^{+0.091}_{-0.063}$ ${}^{+0.057}_{-0.027}$ ${}^{+0.008}_{-0.003}$
$F_{+}$ ${}^{+0}_{-0.02}$ ${}^{+0}_{-0.034}$ ${}^{+0}_{-0.054}$ ${}^{+0.004}_{-0}$ ${}^{+0}_{-0}$
$F_{0}$ ${}^{+0.003}_{-0.023}$ ${}^{+0.007}_{-0.017}$ ${}^{+0.036}_{-0.033}$ ${}^{+0.028}_{-0.025}$ ${}^{+0.008}_{-0}$
$A_{\text{FB,EC}}^{z,z^{\prime}}$ ${}^{+0.05}_{-0.042}$ ${}^{+0}_{-0.1}$ ${}^{+0.078}_{-0.006}$ ${}^{+0.054}_{-0.003}$ ${}^{+0.008}_{-0.041}$
$A_{\text{FB}}^{x^{\prime}}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$
$A_{\text{FB}}^{y^{\prime}}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0.002}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$
$A_{\text{FB}}^{x^{\prime}z^{\prime}}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$ ${}^{+0}_{-0}$
$A_{\text{FB}}^{y^{\prime}z^{\prime}}$ ${}^{+0.008}_{-0}$ ${}^{+0}_{-0.002}$ ${}^{+0.023}_{-0}$ ${}^{+0.01}_{-0}$ ${}^{+0}_{-0}$
Total ${}^{+0.054}_{-0.057}$ ${}^{+0.008}_{-0.11}$ ${}^{+0.13}_{-0.09}$ ${}^{+0.095}_{-0.038}$ ${}^{+0.026}_{-0.041}$
Table 1: Variation in the entanglement measures for different subsystems when the input observables (first column) are allowed to vary within their experimental systematic uncertainty [38, 39, 40]. Entries smaller than $10^{-3}$ are written as zero. For $A_{\text{FB,EC}}^{z,z^{\prime}}$ a measurement is not available, and we assume a systematic uncertainty of ±$\pm 0.1$.
@@ 6 Discussion
In this Letter we have investigated the bipartite and tripartite entanglement between the spins of the top quark decay products $W$, $b$ and their orbital angular momentum $L$. The entanglement could be determined from experimental data already recorded data at the LHC, and without any model-dependent assumptions. The only missing piece of information is a relative phase $e^{i\alpha}$ between amplitudes with $b$ quark helicities ±$\pm 1/2$. As shown, this phase has a very small impact on the entanglement and would hardly influence the interpretation of experimental measurements.
The reader may already have noticed two conspicuous features that arise in our theoretical framework. The $LW$ density operator, obtained after trace on the $b$ spin degrees of freedom, is still sensitive to the phase $e^{i\alpha}$. This is a consequence of the ambiguity in the direct determination of the density operators involving $L$, which we solve by expressing them in terms of $a_{\lambda_{1}\lambda_{2}}$. Even more striking is the fact that the $L$-$b$ entanglement can be established when the $b$ quark spin is not directly measured. This is possible due the change of basis used to write the $LWb$ density operator in terms of a minimal set of parameters $a_{\lambda_{1}\lambda_{2}}$, and the fact that $\lambda=1/2$ amplitudes are suppressed.
In our estimation of systematic uncertainties for entanglement measures we have used ATLAS and CMS analyses that in some cases use small datasets from LHC Run 1. We remark that with higher statistics, as those available with Run 2 data, the full distributions can be used for measurements, not only asymmetries, and the systematic uncertainties are bound to improve. Aiming at a combined measurement of all observables, a deconvolution of the detector effects for measurements of the four-dimensional distribution has also been proposed [41, 42], which also takes into account the correlations among measurements.
In summary, the determination of the bipartite and tripartite entanglement between the spins of top quark decay products and $L$ is based on observables which, except one, have already been measured by the ATLAS or CMS Collaborations, in separate analyses. The current experimental uncertainties suggest that a dedicated analysis of Run 2 data could establish the entanglement between $L$ and the $W$ boson spin, as well as genuine tripartite entanglement, with more than 5 standard deviations. This would be quite a novel type of measurement, which at present has only been achieved with photons.